OpenStudy (anonymous):

solve each system by elimination -1ox-y=-3 -5x+5y=15

OpenStudy (anonymous):

-1x-y=-3 -5x+5y=15 So times first equation by -5: 5x+5y=15 Add both equations to elimeinate x: 10y=30 y=3. Substitute to find x: -x-y=-3 -x-3=-3 -x=0 x=0

OpenStudy (luigi0210):

is that a -10?

OpenStudy (anonymous):

Its -10

OpenStudy (anonymous):

no i hthink he accidentally typed 0

OpenStudy (anonymous):

yeah its -10

OpenStudy (anonymous):

aw man

OpenStudy (anonymous):

i have to redo it now

OpenStudy (anonymous):

-10x-y=-3 -5x+5y=15 So times second equation by -2: 10x-10y=-30 Add both equations to elimeinate x: -11y=-33 y=3. Substitute to find x: -10x-y=-3 -10x-3=-3 -10x=0 x=0. Oh its the same answers still!

OpenStudy (anonymous):

(3, 0)